a current coil of cire 10cm in diameter has 18turns and carries a current of 8a .the coil is inserted into a field of magnetic induction 1.2t what is the maximum torque on the coil?


1
Expert's answer
2021-02-10T12:20:50-0500
"\tau=N I A B\\ \tau=0.25\pi d^2N I B\\\tau=0.25\pi (0.1)^2(18)(8)(1.2)=1.36\ Nm"